Chris is four times as old as Peter. Sara is 6 years younger than Chris. the total age of Chris, Sara and Peter is 66 years. How old is Sara.
Answer by mananth(16946)

let peter's age be x now
chris will be 4x
sara will be4x-6
...
their sum = 66
x+4x+4x-6 =66
9x-6=66
add 6
9x=66+6
9x=72
x= 8
..
sara = 4x-6 = 32-6 = 26 years
...
m.ananth@hotmail.ca